MATTHEW K. GRAY

Term Address 476 Memorial Drive Cambridge, MA 02139

(617) 225–7689 Home Address 207 Thompson Street Port Jefferson, NY 11777

Experience net.Genesis Corporation Boston, MA Founder and Principal Software Architect for an ambitious company offering April, 1994– present

consulting and customization of systems for Internet-based commerce. net.Genesis provides new outlets for sales, advertising and communication between a company and its client using the World Wide Web and other information sources.

O’Reilly and Associates Cambridge, MA Consulting work focusing on producing advertiser reports and performance en-May, 1994– present

hancements to the GNN web server.

MIT MEDIA LAB Cambridge, MA Work on Fishwrap, a project to provide personalized electronic newspapers de-Jan, 1994– present

veloped by the Electronic Publishing group at the Media Lab. Responsibilities included maintenance of servers and other service suppliers, coding in C and perl and helping other members of the project in my areas of expertise.

MIT Radio Astronomy Group Cambridge, MA

January, 1993– Worked on signal processing of light curves attempting to measure time delays May, 1993 in gravitational lense images. Work included C programming and work with

image processing packages. Continuation of earlier work.

Education Massachusetts Institute Of Technology Cambridge, MA Plan to complete B.S. in Physics in June 1995. Classes include computer science and advanced physics classes including programming and lab experience.

Activities  Student Information Processing Board  MIT, Cambridge, MA

Computer services organization at MIT. Chairman from January 1993 to January 1994. Maintain the MIT SIPB World Wide Web Server (http://www.mit.edu:8001/)

Publications “Measuring the Size and Growth of the Worldwide Web,” Global Network Navigator Magazine, January, 1994. (O’Reilly and Associates)